OpenStudy (anonymous):

log a 3≈0.477 and log a(7)≈0.845. Use these values to evaluate log a(147). log a(147)=____

Parth (parthkohli):

\[147 = 7 \cdot 7 \cdot 3 = 7^2 \cdot 3\]Therefore,\[\log(147) \]\[= \log(7^2 \cdot 3) = \log(7^2) + \log(3) = \cdots\]

OpenStudy (cggurumanjunath):